I want to solve the following linear equation.
3(x-7)-2(x-5)=x-11

Answer by fractalier(6550)

Okay, from
3(x-7)-2(x-5)=x-11
let's distribute on the left and get
3x - 21 - 2x + 10 = x - 11
Now collect like terms and get
x - 11 = x - 11
If we subtract off x, we get the curious result
-11 = -11
Well, you have to ask yourself, "When does -11 = -11?"
The answer is "Always!"
Thus, this problem has an infinite number of solutions, for whatever value of x you choose, it works!